OpenStudy (anonymous):

Which expression is equivalent to 3^2 ⋅ 3^−5? fraction 1 over 3 to the power 3 fraction 1 over 3 to the power 7 fraction 1 over 3 to the power negative 3 fraction 1 over 3 to the power negative 7

Nnesha (nnesha):

when multiply same bases you should add their exponents \[\huge\rm x^m \times x^n=x^{m+n}\]

Nnesha (nnesha):

negative exponent \[\huge\rm x^{-m}=\frac{ 1 }{ x^m}\] *flip the fraction*

Nnesha (nnesha):

any question ? did u understand ? yes or no ?

OpenStudy (anonymous):

yeah thanks

OpenStudy (anonymous):

so it will be fraction 1 over 3 to the power negative 7 @Nnesha ?

Nnesha (nnesha):

hmm no first add both exponents and then flip the fraction

OpenStudy (anonymous):

oh so its fraction 1 over 3 to the power 7

Nnesha (nnesha):

\[\huge\rm 3^2 \times 3^{-5} = 3^{??}\]

OpenStudy (anonymous):

i got 0.037... when i multiplied them

Nnesha (nnesha):

well don't use calculator read my first comment here is an example \[\huge\rm 2^5 \times 2^{-6}= 2^{5+(-6)}\]

OpenStudy (anonymous):

its A

Nnesha (nnesha):

how ? explain plz

OpenStudy (anonymous):

i don't know

Nnesha (nnesha):

hmm why did you say its A ? anyways \[\huge\rm 3^2 \times 3^{-5} = 3^{2+(-5)}= 3^{??}\]

Nnesha (nnesha):

when we multiply same bases we should add their exponents